OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by two.

OpenStudy (yanasidlinskiy):

Two I am thinking of are these: (1) pi/5 + 2 pi = pi/5 + 10 pi / 5 = (11 pi )/ 5 and (2) pi/5 - 2 pi = ?

OpenStudy (anonymous):

these are the answers to choose from three pi divided by two; negative pi divided by two pi divided by two + 360°; pi divided by two - 360° five pi divided by two; negative pi divided by two five pi divided by two; negative three pi divided by two

OpenStudy (anonymous):

wait i'm confused, none of the answers have 10pi/5 in them?

OpenStudy (yanasidlinskiy):

Ok. To keep you going...I would recommend you asking @hartnn

OpenStudy (anonymous):

@hartnn any help?

hartnn (hartnn):

to get co-terminal angles add 360 to that angle and subtract 360 from that angle

hartnn (hartnn):

or in radians, add and subtract 2pi

hartnn (hartnn):

so calculate pi/2 + 2pi =... pi/2 - 2pi =...

hartnn (hartnn):

looking at the choice, you don't need to actually calculate them! there is one choice which says to add and subtract 360 degrees :)

OpenStudy (anonymous):

oooh ok got it! thank you!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

@hartnn any help?

hartnn (hartnn):

you still need help with this ? or some new question ?

OpenStudy (anonymous):

Don't put that answer, it is wrong.